OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 30°?

OpenStudy (cwrw238):

sum of exterior angles - 360 degrees so no. of sides = 360/30 = 12 sum of interior angles = 180( 12-2)

OpenStudy (cwrw238):

formula for total angles in polygon = 180(n - 2) degrees

OpenStudy (anonymous):

im lost?

OpenStudy (cwrw238):

you can work out 180(12 - 2)?

OpenStudy (cwrw238):

work out the parentheses first : 12 - 2 = 10

OpenStudy (anonymous):

what is the value of x?

OpenStudy (cwrw238):

x is one exterior angle there are 6 altogether: